Calculate the length of a constant wire of cross sectional areas 4π x 10^8m and resistivity 1.1 x 10^-6Ωm required to construct a standard resistor of resistance 21Ω (take π=22/7)
1
Expert's answer
2020-05-28T11:50:54-0400

The resistance in terms of conductor length "L", area "A" , and resistivity "\rho" is


"R=\frac{\rho L}{A}\rightarrow \\
\space\\L=\frac{RA}{\rho}=\frac{21\cdot4π \cdot10^8 }{1.1 \cdot 10^{-6}}=2.4\cdot10^{16}\text{ m}."

Longer than the distance from the Sun to Saturn!